# Finiteness of totally geodesic hypersurfaces

### v2

@@ 3.4. Proof of Theorem 4
It remains to see how we can restrict the transitivity group so as to argue as above. If $\dim(M)$ is odd, then the argument is exactly as in § 2.2.14. If the dimension is even and the curvature is pinched as in [CLMS23, Theorem 1.2.] then the proof of that theorem shows $H=SO(n-1)$. If $n$ is even and not divisible by $4$ and so equals $24k+$ 2 then following the proof of [CLMS23, Theorem 3.8], we see that unless $n=134$, we have $H<SO(n-2)$. By [Oza91, Theorem 3] we have that $H$ can only be $SO(n-2),U(n-2)$ or $SU(n-2)$ unless $n=10$ or $n=18$. In all such cases we have that $C$ is $[-1,1]$ and we can argue as above.
In dimension $10$ we can in addition have $H=Spin(7)<SO(8)$. By work of Adams, the $9$ sphere only admits one non-vanishing vector field and so the resulting $8$ dimensional representation of $Spin(7)$ cannot have invariant vectors. Therefore the representation is the spin representation of $Spin(7)$ [Ada62]. There is a maximal $G_{2}$ in $Spin(7)$ for which this representation splits off a single trivial representation realizing $Spin(7)/G_{2}$ as $S^{7}$, the unit tangent sphere in our $8$ dimensional representation see e.g. the discussion of maximal subgroups in [Bou68] and of branching rules in [MP81]. It follows from this that $C$ is once again $[-1,1]$ and we argue as before.
In dimension $18$, we can in addition have $H=Spin(9)<SO(16)$. As before, by work of Adams, this is an irreducible representation and so the spin representation of $Spin(9)$. The restriction of this representation to $Spin(8)$ is the sum of the two semi-spin representations of $Spin(8)$. In $Spin(8)$ there are three realizations of $Spin(7)$ coming from triality and for two of these copies of $Spin(7)$ the $16$ dimensional spin representation of $Spin(9)$ splits as a trivial representation, and two irreducibles of dimension $7$ and $8$. See once again [Bou68, MP81]. This shows that there is a $Spin(7)$ invariant vector in the $16$ dimensional representation of $Spin(9)$ and realizes the orbit of that vector as $S^{15}$. This is once again exactly what we need to see that $C$ is $[-1,1]$.
Finally if we are in even dimensions and there is also one closed immersed totally geodesic submanifold $N^{\prime}$ in $M$ then we once again see that the transitivity group is either $SO(n-1)$ or $SO(n-2)$ and the proof works as before.
@@ 4.1.1. Examples in positive curvature
where $h\colon(-1-\varepsilon,1+\varepsilon)\to\mathbb{R}$ is an even, real-analytic function that vanishes at $-1,1$. Note that $h\equiv 0$ yields the standard warped product formula for the round metric, and for $h$ small enough (in the $C^{2}$-sense) the metric will still have positive sectional curvature.
@@ 4.1.2. Examples in negative curvature
The next examples show that there is no local obstruction to a simply connected manifold of non-constant negative curvature having infinitely many totally geodesic hypersurfaces. First, there are many examples of homogeneous variably negatively curved Riemannian manifolds that have infinitely many totally geodesic hypersurfaces. These Riemannian manifolds are isometric to solvable Lie groups equipped with left invariant Riemannian metrics [Hei74]. Lin–Schmidt gave examples of simply connected non-homogeneous negatively curved 3-manifolds each tangent vector to which is contained in a totally geodesic surface [LS17]. In fact, these totally geodesic surfaces can be taken to be isometric to the hyperbolic plane.
@@
for $h$ an even real analytic function as above, $\varphi:\mathbb{R}\rightarrow\mathbb{R}^{\geq 0}$ a smooth monotone non-decreasing function equal to 00 in some neighborhood of 00 and $1$ for all large enough $r$, and $g_{h,\mathbb{S}^{n}}$ the metric on $\mathbb{S}^{n}$ defined in (missing) 4.1.1. For $h\equiv 0$ the $\theta$ variable corresponds to the angular coordinate, and $r$ is equal to the distance to the origin. For any fixed $R>0$, $g_{h,\mathbb{H}^{n+1}}$ can be chosen to be as $C^{2}$-close as desired to the hyperbolic metric for $r<R$, provided that $h$ was chosen with sufficiently small $C^{2}$-norm and $\varphi$ was chosen to have small enough first and second derivatives. The isometries of $g_{\varphi(r)h,\mathbb{S}^{n}}$ described above extend to isometries of $g_{h,\mathbb{H}^{n+1}}$ the fixed point sets of which are totally geodesic hypersurfaces.
Furthermore, provided $g_{h,\mathbb{S}^{n}}$ is close enough to the round metric and $\varphi$ is chosen with small enough first and second derivatives, $g_{h,\mathbb{H}^{n+1}}$ will have negative curvature. One can check this as follows. For points with $r$-coordinate larger than some $R$ depending only on the $C^{2}$ norm of $h$ and $\varphi$, all of their tangent planes will have negative sectional curvature by the formulas that compute the curvature of warped product metrics see [FJ89, Lemma 3.5] and compare [Low21, Lemma 6.2]. The remaining points with $r$ coordinate less than $R$ will then also have negative curvature, provided $g_{h,\mathbb{S}^{n}}$ was chosen sufficiently close to the round metric and $\varphi$ was chosen with sufficiently small first and second derivative, as observed in the previous paragraph.
@@ 4.1.3. Closed negatively curved examples with finitely many totally geodesic hypersurfaces
It is possible to construct smooth Riemannian metrics of negative curvature on a closed hyperbolizable $n$-manifold, $n>2$, that are arbitrarily $C^{2}$-close to the hyperbolic metric and that contain an arbitrarily large number of totally geodesic hypersurfaces. This can be done provided that there are infinitely many closed totally geodesic hypersurfaces in the hyperbolic metric. One constructs such metrics $g_{\epsilon}$ by choosing a finite union of closed totally geodesic hypersurfaces in the hyperbolic (constant sectional curvature -1) metric, and then perturbing the metric on a small ball disjoint from the hypersurfaces in the finite union. We remark that it is sometimes even possible to construct such examples while keeping the sectional curvature bounded above by $-1$ and the totally geodesic hypersurfaces hyperbolic (constant sectional curvature -1 in their induced metric), see [Low23, Section 6].
We can construct analytic metrics $g_{\epsilon}$ as in the previous paragraph as follows. Suppose that $M$ contains infinitely many totally geodesic hypersurfaces in its hyperbolic metric each of which lifts to a finite cover where it is the fixed point set of a finite group of isometries of that finite cover. All first type arithmetic hyperbolic manifolds have infinitely many totally geodesic hypersurfaces of this kind, see [BBKS21]. If we want $k$ totally geodesic hypersurfaces, we choose $g_{\epsilon}$ as above by modifying $g$ away from the $k$ totally geodesic hypersurfaces. We then run Ricci flow for a short time on the metrics $g_{\epsilon}$ of the previous paragraph to obtain metrics $g_{\epsilon}(t)$.
Note that $g_{\epsilon}(t)$ is analytic for $t>0$ by the main result of [Ban87]. Also, since the Ricci flow preserves isometries, and since the fixed point set of a group of isometries is totally geodesic, by pulling back the Ricci flow to the finite covers described in the previous paragraph one can check that each of the $k$ totally geodesic hypersurfaces in $g_{\epsilon}$ is homotopic to a nearby totally geodesic hypersurface in $g_{\epsilon}(t)$.
@@
For one of the metrics $g_{\epsilon}$ as above, we can perturb $g_{\epsilon}$ away from a small neighborhood $U$ of the k totally geodesic hypersurfaces so that no totally geodesic hypersurface in $(M,g_{\epsilon})$ passes through a point of $M-U$. This is possible by using the deformations considered by ([MW19], [EHW24]), which show that a condition on the curvature tensor at point, that is open on the set of Riemannian metrics in the $C^{2}$-topology, prevents a totally geodesic hypersurface from passing through that point.
Call this perturbation $g_{\epsilon}^{\prime}$. Then, running Ricci flow on $g_{\epsilon}^{\prime}$ for a short time, we obtain metrics $g_{\epsilon}^{\prime}(t)$ that have $k$ totally geodesic hypersurfaces homotopic to the $\Sigma_{1},..,\Sigma_{k}$ as above. If we chose $t$ small enough, it will be the case that no point of $M-U^{\prime}$ has a totally geodesic hypersurface in the metric $g_{\epsilon}^{\prime}(t)$ passing through it, for $U^{\prime}$ a slightly larger open neighborhood than $U$ of the union of the $\Sigma_{i}$. Here we are using the fact that the condition on the curvature tensor that ([MW19], [EHW24]) used to check that the perturbed metrics they constructed contained no totally geodesic hypersurfaces was open in the $C^{2}$-topology on the space of Riemannian metrics. Thus, any closed totally geodesic hypersurface $\Sigma$ in $(M,g_{\epsilon}^{\prime}(t))$ must be contained in $U^{\prime}$. Each tangent hyperplane to $\Sigma$ must then be at a distance of at most $\delta$ in the Grassman bundle $Gr_{n-1}(M,g_{\epsilon}(t)^{\prime})$ from a tangent hyperplane to one of the $\Sigma_{i}$, where $\delta$ can be made as small as desired making $U$, $U^{\prime}$, and $t$ small enough. If $\delta$ is chosen small enough, one can then show that normal projection from $\Sigma$ to one of the $\Sigma_{i}$ defines a covering map.This shows that $(M,g_{\epsilon}^{\prime}(t))$ contains exactly $k$ closed maximal totally geodesic hypersurfaces. ∎
It seems possible that one could also do a similar construction in the context of the examples of Gromov and Thurston [GT87] and build metrics with at least $k$ totally geodesic submanifolds on negatively curved manifolds which do not admit metrics of constant negative curvature.
@@ 4.2. Directions for Future Work and Conjectures
As in [BFMS21, BFMS23] given a manifold $M$, we say that an immersed totally geodesic submanifold $N$ in $M$ is maximal if it is not contained in another proper closed immersed totally geodesic submanifold. It is then natural to conjecture the following generalization of Theorem 1:
@@
If the conjecture is true, it then follows from the results of [BFMS21, BFMS23, Cor92, GS92] that $M$ is arithmetic as in Theorem 1. We do not specify the regularity in the conjecture, it seems possible that it holds as soon as the metric is $C^{2}$ and so all the notions in the conjecture are well defined.
The work in [BFMS21, BFMS23] was motivated in part by the Margulis commensurator arithmeticity theorem. In that context, there is an analogy between § 4.2 and results by Eberlein and Farb–Weinberger [Ebe82, FW08]. In particular, the results of Farb–Weinberger suggest there might be an analogue of § 4.2 in the broader context of Riemannian metrics on closed aspherical manifolds.
There are other more general questions than § 4.2 if one thinks of Theorem 1 as a statement about closed manifolds invariant under Anosov flows. It seems possible that an Anosov flow admitting infinitely many closed, flow invariant $C^{1}$ submanifolds is necessarily algebraically defined. Related questions have been considered by Zeghib [Zeg95].

### v3

@@ 3.4. Proof of Theorem 4
It remains to see how we can restrict the transitivity group so as to argue as above. If $\dim(M)$ is odd, then the argument is exactly as in § 2.2.14. If the dimension is even and the curvature is pinched as in [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds, Theorem 1.2.] then the proof of that theorem shows H=SO(n1)H=SO(n-1). If $n$ is even and not divisible by $4$ and so equals $4k+2$ then following the proof of [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds, Theorem 3.8], we see that unless $n=134$, we have H<SO(n2)H<SO(n-2). By [Ozaki, Theorem 3] we have that $H$ can only be SO(n2),U(n2)SO(n-2),U(n-2) or SU(n2)SU(n-2) unless $n=10$ or $n=18$. In all such cases we have that $C$ is [1,1][-1,1] and we can argue as above.
In dimension $10$ we can in addition have $H=Spin(7)<SO(8)$. By work of Adams, the $9$ sphere only admits one non-vanishing vector field and so the resulting $8$ dimensional representation of $Spin(7)$ cannot have invariant vectors. Therefore the representation is the spin representation of $Spin(7)$ [Adams]. There is a maximal $G_{2}$ in $Spin(7)$ for which this representation splits off a single trivial representation realizing $Spin(7)/G_{2}$ as $S^{7}$, the unit tangent sphere in our $8$ dimensional representation see e.g. the discussion of maximal subgroups in [Bourbaki] and of branching rules in [MckayPatera]. It follows from this that $C$ is once again [1,1][-1,1] and we argue as before.
In dimension $18$, we can in addition have $H=Spin(9)<SO(16)$. As before, by work of Adams, this is an irreducible representation and so the spin representation of $Spin(9)$. The restriction of this representation to $Spin(8)$ is the sum of the two semi-spin representations of $Spin(8)$. In $Spin(8)$ there are three realizations of $Spin(7)$ coming from triality and for two of these copies of $Spin(7)$ the $16$ dimensional spin representation of $Spin(9)$ splits as a trivial representation, and two irreducibles of dimension $7$ and $8$. See once again [Bourbaki, MckayPatera]. This shows that there is a $Spin(7)$ invariant vector in the $16$ dimensional representation of $Spin(9)$ and realizes the orbit of that vector as $S^{15}$. This is once again exactly what we need to see that $C$ is [1,1][-1,1].
Finally if we are in even dimensions and there is also one closed immersed totally geodesic submanifold $N^{\prime}$ in $M$ then we once again see that the transitivity group is either SO(n1)SO(n-1) or SO(n2)SO(n-2) and the proof works as before.
@@ 4.1.1. Examples in positive curvature
where $h\colon(-1-\varepsilon,1+\varepsilon)\to\mathbb{R}$ is an even, real-analytic function that vanishes at 1,1-1,1. Note that $h\equiv 0$ yields the standard warped product formula for the round metric, and for $h$ small enough (in the $C^{2}$-sense) the metric will still have positive sectional curvature.
@@ 4.1.2. Examples in negative curvature
The next examples show that there is no local obstruction to a simply connected manifold of non-constant negative curvature having infinitely many totally geodesic hypersurfaces. First, there are many examples of homogeneous variably negatively curved Riemannian manifolds that have infinitely many totally geodesic hypersurfaces. These Riemannian manifolds are isometric to solvable Lie groups equipped with left invariant Riemannian metrics [heintze1974homogeneous]. Lin–Schmidt gave examples of simply connected non-homogeneous negatively curved 3-manifolds each tangent vector to which is contained in a totally geodesic surface [ls17]. In fact, these totally geodesic surfaces can be taken to be isometric to the hyperbolic plane.
@@
for $h$ an even real analytic function as above, $\varphi:\mathbb{R}\rightarrow\mathbb{R}^{\geq 0}$ a smooth monotone non-decreasing function equal to 0 in some neighborhood of 0 and $1$ for all large enough $r$, and $g_{h,\mathbb{S}^{n}}$ the metric on $\mathbb{S}^{n}$ defined in () 4.1.1. For $h\equiv 0$ the $\theta$ variable corresponds to the angular coordinate, and $r$ is equal to the distance to the origin. For any fixed $R>0$, $g_{h,\mathbb{H}^{n+1}}$ can be chosen to be as $C^{2}$-close as desired to the hyperbolic metric for $r<R$, provided that $h$ was chosen with sufficiently small $C^{2}$-norm and $\varphi$ was chosen to have small enough first and second derivatives. The isometries of $g_{\varphi(r)h,\mathbb{S}^{n}}$ described above extend to isometries of $g_{h,\mathbb{H}^{n+1}}$ the fixed point sets of which are totally geodesic hypersurfaces.
Furthermore, provided $g_{h,\mathbb{S}^{n}}$ is close enough to the round metric and $\varphi$ is chosen with small enough first and second derivatives, $g_{h,\mathbb{H}^{n+1}}$ will have negative curvature. One can check this as follows. For points with $r$-coordinate larger than some $R$ depending only on the $C^{2}$ norm of $h$ and $\varphi$, all of their tangent planes will have negative sectional curvature by the formulas that compute the curvature of warped product metrics see [FJ, Lemma 3.5] and compare [l21, Lemma 6.2]. The remaining points with $r$ coordinate less than $R$ will then also have negative curvature, provided $g_{h,\mathbb{S}^{n}}$ was chosen sufficiently close to the round metric and $\varphi$ was chosen with sufficiently small first and second derivative, as observed in the previous paragraph.
@@ 4.1.3. Closed negatively curved examples with finitely many totally geodesic hypersurfaces
It is possible to construct smooth Riemannian metrics of negative curvature on a closed hyperbolizable $n$-manifold, $n>2$, that are arbitrarily $C^{2}$-close to the hyperbolic metric and that contain an arbitrarily large number of totally geodesic hypersurfaces. This can be done provided that there are infinitely many closed totally geodesic hypersurfaces in the hyperbolic metric. One constructs such metrics $g_{\epsilon}$ by choosing a finite union of closed totally geodesic hypersurfaces in the hyperbolic (constant sectional curvature -1) metric, and then perturbing the metric on a small ball disjoint from the hypersurfaces in the finite union. We remark that it is sometimes even possible to construct such examples while keeping the sectional curvature bounded above by 1-1 and the totally geodesic hypersurfaces hyperbolic (constant sectional curvature -1 in their induced metric), see [Lowe, Section 6].
We can construct analytic metrics $g_{\epsilon}$ as in the previous paragraph as follows. Suppose that $M$ contains infinitely many totally geodesic hypersurfaces in its hyperbolic metric each of which lifts to a finite cover where it is the fixed point set of a finite group of isometries of that finite cover. All first type arithmetic hyperbolic manifolds have infinitely many totally geodesic hypersurfaces of this kind, see [belolipetsky2021subspace]. If we want $k$ totally geodesic hypersurfaces, we choose $g_{\epsilon}$ as above by modifying $g$ away from the $k$ totally geodesic hypersurfaces. We then run Ricci flow for a short time on the metrics $g_{\epsilon}$ of the previous paragraph to obtain metrics $g_{\epsilon}(t)$.
Note that $g_{\epsilon}(t)$ is analytic for $t>0$ by the main result of [Bando]. Also, since the Ricci flow preserves isometries, and since the fixed point set of a group of isometries is totally geodesic, by pulling back the Ricci flow to the finite covers described in the previous paragraph one can check that each of the $k$ totally geodesic hypersurfaces in $g_{\epsilon}$ is homotopic to a nearby totally geodesic hypersurface in $g_{\epsilon}(t)$.
@@
For one of the metrics $g_{\epsilon}$ as above, we can perturb $g_{\epsilon}$ away from a small neighborhood $U$ of the k totally geodesic hypersurfaces so that no totally geodesic hypersurface in $(M,g_{\epsilon})$ passes through a point of MUM-U. This is possible by using the deformations considered by ([MurphyWilhelm], [ElHWilhelm]), which show that a condition on the curvature tensor at point, that is open on the set of Riemannian metrics in the $C^{2}$-topology, prevents a totally geodesic hypersurface from passing through that point.
Call this perturbation $g_{\epsilon}^{\prime}$. Then, running Ricci flow on $g_{\epsilon}^{\prime}$ for a short time, we obtain metrics $g_{\epsilon}^{\prime}(t)$ that have $k$ totally geodesic hypersurfaces homotopic to the $\Sigma_{1},..,\Sigma_{k}$ as above. If we chose $t$ small enough, it will be the case that no point of $M-U^{\prime}$ has a totally geodesic hypersurface in the metric $g_{\epsilon}^{\prime}(t)$ passing through it, for $U^{\prime}$ a slightly larger open neighborhood than $U$ of the union of the $\Sigma_{i}$. Here we are using the fact that the condition on the curvature tensor that ([MurphyWilhelm], [ElHWilhelm]) used to check that the perturbed metrics they constructed contained no totally geodesic hypersurfaces was open in the $C^{2}$-topology on the space of Riemannian metrics. Thus, any closed totally geodesic hypersurface $\Sigma$ in $(M,g_{\epsilon}^{\prime}(t))$ must be contained in $U^{\prime}$. Each tangent hyperplane to $\Sigma$ must then be at a distance of at most $\delta$ in the Grassman bundle $Gr_{n-1}(M,g_{\epsilon}(t)^{\prime})$ from a tangent hyperplane to one of the $\Sigma_{i}$, where $\delta$ can be made as small as desired making $U$, $U^{\prime}$, and $t$ small enough. If $\delta$ is chosen small enough, one can then show that normal projection from $\Sigma$ to one of the $\Sigma_{i}$ defines a covering map.This shows that $(M,g_{\epsilon}^{\prime}(t))$ contains exactly $k$ closed maximal totally geodesic hypersurfaces. ∎
It seems possible that one could also do a similar construction in the context of the examples of Gromov and Thurston [GromovThurston] and build metrics with at least $k$ totally geodesic submanifolds on negatively curved manifolds which do not admit metrics of constant negative curvature.
@@ 4.2. Directions for Future Work and Conjectures
As in [BFMS1, BFMS2] given a manifold $M$, we say that an immersed totally geodesic submanifold $N$ in $M$ is maximal if it is not contained in another proper closed immersed totally geodesic submanifold. It is then natural to conjecture the following generalization of Theorem 1:
@@
If the conjecture is true, it then follows from the results of [BFMS1, BFMS2, Corlette, GromovSchoen] that $M$ is arithmetic as in Theorem 1. We do not specify the regularity in the conjecture, it seems possible that it holds as soon as the metric is $C^{2}$ and so all the notions in the conjecture are well defined.
The work in [BFMS1, BFMS2] was motivated in part by the Margulis commensurator arithmeticity theorem. In that context, there is an analogy between § 4.2 and results by Eberlein and Farb–Weinberger [Eberlein1, FarbWeinberger]. In particular, the results of Farb–Weinberger suggest there might be an analogue of § 4.2 in the broader context of Riemannian metrics on closed aspherical manifolds.
There are other more general questions than § 4.2 if one thinks of Theorem 1 as a statement about closed manifolds invariant under Anosov flows. It seems possible that an Anosov flow admitting infinitely many closed, flow invariant $C^{1}$ submanifolds is necessarily algebraically defined. Related questions have been considered by Zeghib [Zeghib1995_Sur-une-notion-dautonomie-de-systemes-dynamiques-appliquee-aux-ensembles-invariants].